# Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents

### v1

@@ Abstract
Uncertainty quantification (UQ) for large language models (LLMs) is a key building block for safety guardrails of daily LLM applications. Yet, even as LLM agents are increasingly deployed in highly complex tasks, most UQ research still centers on single-turn question-answering. We argue that UQ research must shift to realistic settings with interactive agents, and that a new principled framework for agent UQ is needed. This paper presents the first general formulation of agent UQ that subsumes broad classes of existing UQ setups. Under this formulation, we show that prior works implicitly treat LLM UQ as an uncertainty accumulation process, a viewpoint that breaks down for interactive agents in an open world. In contrast, we propose a novel perspective, a conditional uncertainty reduction process, that explicitly models reducible uncertainty over an agent’s trajectory by highlighting “interactivity” of actions. From this perspective, we outline a conceptual framework to provide actionable guidance for designing UQ in LLM agent setups. Finally, we conclude with practical implications of the agent UQ in frontier LLM development and domain-specific applications, as well as open remaining problems.
@@ 1 Introduction
Figure 1: Comparison between UQ setups. Traditional LLM UQ (a) measures the uncertainty of an answer given a question, whereas the UQ for LLM reasoning (b) expands the problem by considering multi-step responses rather than a single response. Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau$-bench Airline; barres2025tau).
@@
LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent] at the next level of autonomy [Steinberger2026OpenClaw]. In such settings, failures are no longer limited to incorrect text generation: agents may act prematurely under unresolved ambiguity, propagate errors across long interaction trajectories, or commit to outcomes that are costly or difficult to undo. To be deployed responsibly, agents must be able to assess and act upon the likelihood of failure [kochenderfer2015decision]. This makes reliable uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
@@
Despite the importance, most existing UQ research treats LLMs as static oracles: a system examined in isolation, prompted once, and evaluated by the uncertainty of a single response or a chain of responses (Figure 1, left). These methods implicitly assume a static system in which no new information is acquired after the initial prompt [malinin2021uncertainty, kadavath2022language, farquhar2024detecting]. Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation. While appropriate for single-turn inference and non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
@@
Interactive agents operate in multi-turn interactive environments with feedback from users, tools, and external systems (Figure 1, right). For example, a flight-booking agent may need to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences. Early uncertainty should prompt information-seeking actions, while subsequent dialogue and database queries can resolve ambiguities and reduce uncertainty, enabling confident execution. Here, uncertainty does not strictly accumulate–it may also decrease as a direct consequence of interaction, even as the trajectory grows longer.
In this paper, we argue that making LLM agents trustworthy requires a shift in focus: from pointwise uncertainty of answers to structured uncertainty flow in open, interactive decision processes; and a new framework is necessary for this. We present the following perspective:
Uncertainty quantification for an LLM agent must incorporate interactivity to model reducible uncertainty: in non-isolated, agentic systems, total uncertainty should be allowed to decrease as new information is acquired through interaction.
@@
To ground this perspective in a formal discussion, we first provide a concrete definition of agent UQ and a general formulation that captures broad categories of existing UQ setups (Sec. 3). Specifically, we model the agent’s problem-solving trajectory as a stochastic process over actions, observations, and environment states, represented as a simple graphical model. Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ approaches arise as special cases.
This unified view also exposes a fundamental limitation of existing methods: uncertainty is implicitly treated as accumulating monotonically along the trajectory, regardless of the role played by different actions; thus fails to relate uncertainty to a task reward. To address this limitation, we introduce in Sec. 4 a conceptual framework that models agent UQ as a conditional uncertainty reduction process, through an information-gating mechanism that operationalizes the distinction between uncertainty-increasing and uncertainty-reducing actions by considering their interactivity. We further derive the analytic bounds that improve the interpretability of the resulting uncertainty estimates.
@@
In addition, Sec. 5 discusses practical implications of agent UQ across diverse domains, including healthcare, software engineering, and robotics, as well as frontier LLM research on adaptive reasoning and post-training. Finally, we outline open problems unique to uncertainty modeling in agentic systems (Sec. 6). Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware agentic systems.
  • We formalize agent UQ as a distinct problem setup and show why existing LLM UQ formulations are insufficient for interactive agent setups.
  • To tackle that limitation, we present a conceptual framework for agent UQ by viewing it as a conditional uncertainty reduction process.
@@
  • To give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
@@ 2 Related Work
In contrast to classic UQ [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs. Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over samples [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal]. While promising, they mostly use single-step question-answering as a testbed, which limits their practical applicability to agentic setups.
Figure 2: Graphical model for an agent problem-solving trajectory with examples. Given a task specification $E_{0}$ and an initial user query $O_{0}$, an agent spans a multi-turn trajectory characterized by a chain of action AA, observation $O$, and environment state $E$. This simple abstraction describes some representative agentic prompting methods such as ReAct [yao2022react]. See Appendix A.1 for details.
@@
Beyond single-step generation, researchers show growing interest in measuring the confidence of LLMs during their reasoning. This kind of (un)certainty information has been used to improve the quality of response [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning]. Recently, there have also been attempts to model uncertainty in interactive and/or agentic setups [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal]. Although UQ here aims to model uncertainty dynamics over long-horizon generation, existing methods do not capture the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new framework for agent UQ.
@@ 3.1 Problem Statement and Definitions
Imagine a flight booking agent depicted in Figure 1 right. Here, the types of action $A$ and observation $O$ are diverse: for $A$, there is information-seeking, questioning for clarification, calling an API function, and thinking or updating without interactive behaviors. The observation $O$, on the other hand, can be a database status report, tool-calling results, or the user’s clarification. Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., code.
@@
Let $E_{i}$ be an environment state, a mixture of a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and a state of the system database, which is not directly accessible to the agent (only partially observable via tool-calling). Let $O_{i}$ and $A_{i}$ be the $i$-th turn observation and action derived from distributions $P$ and $P_{\pi,\mathcal{T}}$, respectively, where $\mathcal{T}$ indicates a tool set and $\pi$ is an LLM policy. Then, under a task specification $E_{0}$ and the user’s initial query $O_{0}$, we define the trajectory as a sequence of turns $\mathcal{F}_{\leq T}=\{(A_{t},E_{t},O_{t})\}_{t=0}^{T}$ where $\mathcal{F}_{0}=(E_{0},O_{0})$. Given a deterministic update function $h(\cdot)$, the generative process for each step $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
To express this trajectory with a simple mathematical representation, we introduce a dynamic Bayesian network [murphy2002dynamic] in Figure 2. With this graphical model, dependency structures of $A$, $E$, and $O$ can be explained as follows: (1) since the environment state $E_{i-1}$ contains the entire conversation history of $A_{\leq i-1}$ and $O_{<i-1}$ up to the i1i-1 step [^1] An adaptive memory structure [xu2025mem, xiong2025memory] can replace this ever-growing memory., the current action $A_{i}$ depends only on $E_{i-1}$ and $O_{i-1}$; (2) the current observation $O_{i}$ from the environment also depends only on $A_{i}$ and $E_{i}$; (3) the transition of the environment state ($E_{i-1}\rightarrow E_{i}$) can be defined as a deterministic function [^2] In general, this system transition can be also stochastic, but we consider a simplified yet still realistic scenario. $h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system state given the current action.
@@
In the machine learning field, the most popular choices of the uncertainty function $U(X)$ are Shannon entropy [shannon1948mathematical] $\mathbb{E}[-\log P(X)]$, negative log probability $\log P(X=x)$, and their variants. These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suit for sequential modeling. For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix A and B for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
@@
$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$
where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ denotes an initial query uncertainty plus a task volatility.
@@
We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus. See Appendix C for discussion.
@@ 3.2 A Unified View on Existing UQ Setups
where $w_{i}\geq 0$ and $\sum_{i=1}^{T}w_{i}=1$. Here, $O_{0}$ and $A_{T}$ denote the given initial query and the model’s final response, and $A_{i},...,A_{T-1}$ are intermediate outputs. Similar to the single-step setup, most existing multi-step UQs [tanneru2024quantifying, tao2025revisiting] do not consider the question ambiguity $U(O_{0})$, so that reduces Eq. 2 to Eq. 3, while a few consider it [leang2025picsar]. Meanwhile, rather than summing up all the step-wise uncertainties, some approaches consider the most uncertain one, i.e., the lowest confidence used in fu2025deep, which can be represented by Eq. 4. Finally, Eq. 5 expresses various weighted average methods, such as a standard length-normalized uncertainty [kang2025scalable] by setting $w_{i}=\frac{1}{T}$, and a tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies, i.e., top-K mean [fu2025deep] or model-based estimates [zhao2025uncertainty].

### v2

# Uncertainty Quantification in LLM Agents: Foundations, Emerging Challenges, and Opportunities

@@ Abstract
Uncertainty quantification (UQ) for large language models (LLMs) is a key building block for safety guardrails of daily LLM applications. Yet, even as LLM agents are increasingly deployed in highly complex tasks, most UQ research still centers on single-turn question-answering. We argue that UQ research must shift to realistic settings with interactive agents, and that a new principled framework for agent UQ is needed. This paper presents three pillars to build a solid ground for future agent UQ research: (1. Foundations) We present the first general formulation of agent UQ that subsumes broad classes of existing UQ setups; (2. Challenges) We identify four technical challenges specifically tied to agentic setups—selection of uncertainty estimator, uncertainty of heterogeneous entities, modeling uncertainty dynamics in interactive systems, and lack of fine-grained benchmarks—with numerical analysis on a real-world agent benchmark, $\tau^{2}$-bench; (3. Future Directions) We conclude with noting on the practical implications of agent UQ and remaining open problems as forward-looking discussion for future explorations.
@@ 1 Introduction
Figure 1: Comparison between UQ setups. Traditional LLM UQ (a) measures the uncertainty of an answer given a question, whereas the UQ for LLM reasoning (b) expands the problem by considering multi-step responses rather than a single response. Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau^{2}$-bench Airline; barres2025tau).
@@
LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent, anthropic2026cowork] at the next level of autonomy [Steinberger2026OpenClaw]. In such settings, failures are no longer limited to incorrect text generation: agents may act prematurely under unresolved ambiguity, propagate errors across long interaction trajectories, or commit to outcomes that are costly or difficult to undo. To be deployed responsibly, agents must be able to assess and act upon the likelihood of failure [kochenderfer2015decision]. This makes uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
@@
Despite the importance, most existing UQ research treats LLMs as static oracles: a system examined in isolation, prompted once, and evaluated by the uncertainty of a single response or a chain of responses (Figure 1, left). These methods implicitly assume a static system in which no new information is acquired after the initial prompt [malinin2021uncertainty, kadavath2022language, farquhar2024detecting]. Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation process. While appropriate for single-turn inference or non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
@@
LLM agents in this setup operate with feedback from users and system components through multi-turn interactions to complete a complex task (Figure 1, right). For example, a flight-booking agent needs to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences. Early uncertainty should prompt information-seeking actions, while subsequent dialogue and database queries can resolve ambiguities and reduce uncertainty, enabling confident execution. This deviates significantly from the setup considered in classic LLM UQ, as the agent now spans a long interaction log updated over multiple turns that involves messages from other entities, and it also gets additional information from environments. This gap raises questions about the applicability of existing UQ methods to agentic setups.
@@
In this paper, we argue that building a trustworthy LLM-driven agentic system requires a shift in focus: from pointwise uncertainty of the final answer to structured uncertainty dynamics in an open, interactive decision process, and a new framework is necessary for this. To initiate a formal discussion, we first provide a concrete definition and a general formulation of agent UQ that captures broad categories of existing UQ setups (Sec. 3). Specifically, we model the agent’s problem-solving trajectory as a stochastic process over actions, observations, and environment states, represented as a simple graphical model. Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ problems can be cast as special cases.
Based on this foundation, we identify four technical challenges in agent UQ that call for actions from communities. (1) Selection of uncertainty estimator: while it is well-known that existing uncertainty estimators have their own weaknesses, agentic scaffolding amplifies these into severe limitations or hard constraints; (2) Uncertainty of heterogeneous entities: how can we derive estimates over the content generated by external entities having different underlying distributions?; (3) Modeling uncertainty dynamics in interactive systems: traditional weighted-average-style uncertainty aggregation may not be suitable to modeling uncertainty dynamics in the non-isolated, open environment; (4) Lack of fine-grained benchmarks: our pilot survey reveals that granular benchmarks, such as turn-level evaluation, are scarce for LLM agents.
@@
Then, Sec. 5 discusses practical implications of agent UQ across diverse domains, including healthcare, software engineering, and robotics, as well as frontier LLM research on adaptive reasoning and post-training. Finally, we outline open problems specific to uncertainty modeling in agentic systems (Sec. 6). Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware, reliable agentic systems.
  • We cast agent UQ as a distinct problem setup, and build a foundation for that by providing a concrete definition and a general formulation that expresses many existing UQ setups as special cases.
  • We crystallize the emerging challenges in agent UQ with numerical studies on a real-world benchmark $\tau^{2}$-bench [barres2025tau] with GPT-4.1 [openai2025gpt41] and Kimi-K2.5 [kimi2026].
@@
  • To further give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
@@ 2 Related Work
In contrast to classic UQ in machine learning [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs. Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over multiple generations [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal]. While promising, they mostly use single-step question-answering as a testbed, which limits their practical applicability to agentic setups.
Figure 2: Graphical model for an agent problem-solving trajectory with examples. Given a task specification $E_{0}$ and an initial user query $O_{0}$, an agent spans a multi-turn trajectory characterized by a chain of action AA, observation $O$, and environment state $E$. This simple abstraction describes some representative agentic prompting methods such as ReAct [yao2022react]. See Appendix B.1 for details.
@@
Beyond single-step generation, researchers show growing interest in measuring the confidence of LLMs during their reasoning. This (un)certainty information has been used to improve response quality [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning]. Recently, there have been attempts to model uncertainty in interactive or agentic inference [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal, zhang2026confidence, zhang2026auq]. Although aiming to model uncertainty dynamics over long-horizon generation, they do not reflect uncertainty from different entities (e.g., users) nor concern the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new UQ framework under agentic scaffolding.
@@ 3.1 Problem Statement and Definitions
Imagine a flight booking agent depicted in Figure 1 right. Here, the types of action $A$ and observation $O$ are diverse: for $A$, there is information-seeking, questioning for clarification, calling an API function, and thinking or updating without interactive behaviors. The observation $O$, on the other hand, can be a database status report, tool-calling results, or the user’s clarification. Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., programming language.
@@
Let $E_{i}$ be an environment state, a mixture of (1) a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and (2) a state of the system database, which is not directly accessible to the agent—only partially observable via tool-calling. Let $O_{i}$ and $A_{i}$ be the $i$-th turn observation and action derived from distributions $P$ and $P_{\pi,\mathcal{T}}$, respectively, where $\mathcal{T}$ indicates a tool set and $\pi$ is an LLM policy. Then, under a task specification $E_{0}$ and the user’s initial query $O_{0}$, we define the trajectory as a sequence of turns $\mathcal{F}_{\leq T}=\{(A_{t},E_{t},O_{t})\}_{t=0}^{T}$ where $\mathcal{F}_{0}=(E_{0},O_{0})$. Given a deterministic update function $h(\cdot)$, the generative process for each turn $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
To express this trajectory with a simple mathematical representation, we introduce a dynamic Bayesian network [murphy2002dynamic] in Figure 2. With this graphical model, dependency structures of $A$, $E$, and $O$ can be explained as follows: (1) since the environment state $E_{i-1}$ contains the entire conversation history of $A_{\leq i-1}$ and $O_{<i-1}$ up to the i1i-1 step [^1] An adaptive memory structure [xu2025mem, xiong2025memory] can replace this ever-growing memory., the current action $A_{i}$ depends only on $E_{i-1}$ and $O_{i-1}$; (2) the current observation $O_{i}$ from the environment also depends only on $A_{i}$ and $E_{i}$; (3) the transition of the environment state ($E_{i-1}\rightarrow E_{i}$) can be defined as a deterministic function [^2] In general, this system transition can be also stochastic, but we consider a simplified yet still realistic scenario. $h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system database state given the current action.
@@
In the machine learning field, the most popular choices of the uncertainty function $U(X)$ are Shannon entropy [shannon1948mathematical] $\mathbb{E}[-\log P(X)]$, negative log probability $\log P(X=x)$, and their variants. These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suitable for sequential modeling. For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix B and C for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
@@
$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$ (1)
where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ is an initial query uncertainty plus a task specification volatility.
@@
We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded by, while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus. See Appendix D for the relevant discussion.
@@ 3.2 A Unified View on Existing UQ Setups
where $w_{i}\geq 0$ and $\sum_{i=1}^{T}w_{i}=1$. Here, $O_{0}$ and $A_{T}$ denote the given initial query and the model’s final response, and $A_{i},...,A_{T-1}$ are intermediate outputs. Similar to the single-step setup, most existing multi-step UQs [tanneru2024quantifying, tao2025revisiting] do not consider the question ambiguity $U(O_{0})$, so that reduces Eq. 3 to Eq. 4, while a few consider it [leang2025picsar]. Meanwhile, rather than summing up all the step-wise uncertainties, some approaches consider the most uncertain one, i.e., the lowest confidence used in fu2025deep, which can be represented by Eq. 5. Finally, Eq. 6 expresses various weighted average methods, e.g., length-normalized uncertainty [kang2025scalable] with $w_{i}=\frac{1}{T}$, and tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies [zhao2025uncertainty].